# Efficient fault-tolerant code switching via one-way transversal CNOT gates

### v1

@@ Abstract
Code switching is an established technique that facilitates a universal set of fault-tolerant (FT) quantum gate operations by combining two quantum error correcting (QEC) codes with complementary sets of gates, which each by themselves are easy to implement fault-tolerantly. In this work, we present a code switching scheme that respects the constraints of FT circuit design by only making use of transversal gates. These gates are intrinsically FT without additional qubit overhead. We analyze application of the scheme to low-distance color codes, which are suitable for operation in existing quantum processors, for instance based on trapped ions or neutral atoms. We also briefly discuss connectivity constraints that arise for architectures based on superconducting qubits. Numerical simulations of circuit-level noise indicate that a logical -gate, facilitated by our scheme, could outperform both flag-FT magic state injection protocols and a physical -gate at low physical error rates. Transversal code switching naturally scales to code pairs of arbitrary code distance. We observe improved performance of a distance-5 protocol compared to both the distance-3 implementation and the physical gate for realistically attainable physical entangling gate error rates. We discuss how the scheme can be implemented with a large degree of parallelization, provided that logical auxiliary qubits can be prepared reliably enough. Our logical -gate circumvents the need for potentially costly magic state factories. The requirements to perform QEC and to achieve an FT universal gate set are then essentially the same: Prepare logical auxiliary qubits offline, execute transversal gates (ideally in parallel) and perform fast-enough measurements. Transversal code switching thus serves to enable more practical hardware realizations of FT universal quantum computation. The scheme alleviates resource requirements for experimental demonstrations of quantum algorithms run on logical qubits.
I Introduction
@@
A prominent modern technique to render quantum circuits FT with little qubit and gate overhead is the flag qubit paradigm [12, 13, 14, 15, 16, 17, 18, 19, 20]. Faults within the circuit that may cause uncorrectable errors on the logical qubit are heralded by controlled propagation via additional CNOT gates, which are interleaved into the quantum circuit in an intricate way, onto extra auxiliary qubits. When a flag qubit is measured in a particular state, one may employ further circuity to treat such faults or repeat the circuit altogether until the flag is clear. On the one hand, flag circuits may be favorable due to their conceptual simplicity and typically small qubit overhead. On the other hand, dynamical circuit branching with relatively long circuit sequences might pose a challenge for practical applications of QEC with flag circuits. Also, recent developments of quantum computing hardware suggest that the number of qubits may no longer be the most critical limiting factor [21, 22, 23, 24, 25].
@@
For the controlled manipulation of encoded information, FT gate operations must be performed on the logical qubits. Transversal CNOT gates are appealing because they do not require a repetition overhead to respect fault tolerance such as, e.g., implementations via lattice surgery [30, 31, 32]. It is known that no non-trivial QEC code offers a transversal universal gate set [33]. Therefore, given any code with a set of transversal gates, additional strategies are imperative to establish universality. There exist several procedures such as gate teleportation [34] or code conversion [35, 36] that vary in intricacy from both a theoretical and experimental point of view. Code switching in particular refers to the idea of employing two different QEC codes, which each on their own have a different set of transversal gates but in conjunction offer an FT universal gate set that is simple to perform on the logical level [37, 38]. The difficulty of implementing a particular gate is traded in for transferring the logical information between the two codes at will in an FT fashion.
@@
In this work, we investigate a new approach to implement the non-Clifford -gate fault-tolerantly on the 2D color code via FT code switching to the 3D color code that is based on code doubling [43, 44, 45, 46]. A notable strength of the scheme is that it can be viewed as deterministic in practice: Only QEC codes of distance are employed, instead of mere quantum error detecting codes with , and no post-selection is required. The scheme is considered efficient because it relies on a transversal construction of the logical CNOT gate, which works as long as one restricts the connection of the gate to be one-directional. Transversal code switching is performed via logical state teleportation. The logical -gate implementation is then built on two switching steps “2D 3D 2D”, depicted with logical building blocks in Fig. 1. Logical auxiliary qubits in both codes can be prepared with Clifford circuits only, so that no complicated magic state subroutines are needed.
This manuscript is organized as follows: The efficient FT -gate construction is discussed in detail in Sec. II. We offer insights on parallelization and embedding into a reduced-connectivity hardware architecture. In Sec. III, we present simulations of incoherent circuit-level Pauli noise with both a simple single-parameter depolarizing noise model and a more elaborate multi-parameter noise model that can be used to capture the essential noise processes of a trapped-ion quantum computer. A comparison to other strategies for establishing an FT universal gate set is provided in Sec. IV. In Sec. V, we consider scaling up our scheme to QEC codes of arbitrary distance and lastly draw conclusions and provide an outlook on future work in Sec. VI.
II Transversal code switching
@@
Essential for any distance- QEC code, FT code switching must ensure that no more than errors of any Pauli type or can result from up to fault events in a circuit-level noise model. Such faults must be assumed to happen during the execution of the actual quantum circuits of the switching algorithm where every physical operation acting on qubits is subjected to a -qubit depolarizing noise channel. A logical gate is deemed transversal if no interactions between physical qubits happen within the same logical block; faults are confined to the qubits on which they initially occur this way and never spread uncontrollably.
@@
II.1 Doubled-code construction
@@
In Ref. [46], a code construction is described that allows one to perform FT code switching with the help of logical auxiliary qubits and one-way transversal CNOT gates. The central ingredients are a self-dual Calderbank-Shor-Steane (CSS) code111This need not be a color code. and a triorthogonal code, from which a so-called doubled code is constructed, which is itself again a triorthogonal code with parameters . The stabilizers and logical operators of the two ingredient codes are recombined to a new, bigger triorthogonal matrix in the process of code doubling. The initial self-dual code is complemented with additional -type stabilizers to yield the new triorthogonal code, which may have non-local stabilizers. Importantly, the original -type stabilizers of the self-dual code remain intact and are incorporated into the doubled code. As a consequence, the one-way logical CNOT gate can be executed by transversally connecting physical CNOT gates between the physical qubits of the self-dual code and a subset of physical qubits of the doubled code. This enables bi-directional switching between the two codes via standard quantum state teleportation circuits at the logical level. Transversality ensures that fault tolerance is respected for a code switching step. Additionally, it must be guaranteed that no uncorrectable error can ever propagate between the two codes from the preparation of the logical auxiliary qubits. While code doubling has been investigated before [43, 44, 45], the one-way transversal CNOT gate is a new ingredient identified in Ref. [46].
II.2 Reduction to 2D/3D color code combination
@@
The logical CNOT gate can be performed directly between a 2D and a 3D color code at the price of restricting it to only one direction. The reason for this is that the 3D color code already is a doubled code. For the code definitions used in Ref. [46], this means that the control of the logical CNOT may only be encoded in the 3D color code and the target of the logical CNOT must be encoded in the 2D color code. Using only physical CNOT gates, it can be ensured that all stabilizer operators of each code propagate through the logical CNOT onto stabilizer operators of the respective other code. The product of the stabilizer groups of both codes is therefore left invariant. Simultaneously, logical operators of both - and -type are mapped between the two codes as prescribed by the action of a logical CNOT gate.
@@
The remainder of this section is dedicated to applying the aforementioned code switching protocol to the smallest instances of the 2D and 3D color code respectively depicted in Fig. 2, i.e., the seven-qubit Steane code and the fifteen-qubit Tetrahedral code. Both QEC codes have distance and thus correct arbitrary Pauli error. FT state preparation of the logical auxiliary qubit states will be discussed in Sec. II.2.2. After providing a detailed discussion of this small-scale implementation, we broaden the horizon again and discuss scaling up to large distances in Sec. V.
II.2.1 Distance-3 transversal code switching
@@
We now also consider the propagation of errors. It is obvious that a single correctable error in the Steane code can at most propagate to a single error on the Tetrahedral code, which is also correctable, due to transversality. Correctable -errors of weight-3 in the Tetrahedral code may either propagate to correctable weight-1 errors in the Steane code through a single physical CNOT gate but can also, and more severely, directly propagate to uncorrectable weight-3 -errors on the Steane code, as shown in Fig. 4b. Their appearance must therefore be prevented by an appropriate FT state preparation circuit. It is not enough to only fault-tolerantly prepare the 3D color code state but the state preparation routine must be FT with respect to the 2D color code as well in order to render the whole code switching scheme FT.
@@ II.2.2 FT state preparation
2) High-weight -errors: -errors of weight larger than 1 on must be prevented from propagating to the Steane code during the course of code switching although they could be correctable in the Tetrahedral code. It is clear from Figs. 1 and 4b that propagation of such an error to the Steane code via the first logical CNOT gate, may lead to a false decoding result and subsequent erroneous application of a logical -operator. We find that, in our state preparation subroutine, three flags are necessary to catch all high-weight -errors. These flags can be chosen to measure the stabilizers , and .
@@
For practical purposes, we suggest to use these circuits for non-deterministic logical state preparation because the acceptance rate of approximately 333This value is estimated from a circuit-level depolarizing noise model with parameters and . More detail will be provided in Sec. III. is reasonably close to unity for state-of-the-art quantum processors where the entangling gate error rate of order is the limiting factor [52]. No mid-circuit measurements or branching logic are required to correctly flag all dangerous faults. Deterministic variants of logical state preparation might as well be employed but we suspect this approach to require a larger number of entangling gates in general for low physical error rates.
@@ II.3 Parallelized implementation
Given logical auxiliary qubits and parallelized physical-level gates, the transversal code switching scheme can be implemented in constant circuit depth w.r.t. the code size. The first logical CNOT requires one time step, the transversal measurements and corrections require one time step each, the physical -gates are all applied in parallel and the second code switching step then is done again in three time steps; the full logical -gate circuitry is thus performed in seven steps regardless of the code distances, assuming a supply of logical auxiliary qubits is available. Unitary encoding circuits for the preparation of -qubit logical auxiliary states with depth can be found for 2D topological codes [59] and might be improved by exploiting symmetries [60] or with the help of physical auxiliary qubits [61].
@@ II.4.1 Superconducting architecture
The long-range connectivity required for unitary encoding circuits could in theory be circumvented with the help of circuit knitting [64]. One may resort to a stabilizer-measurement-based state preparation routine altogether if non-destructive operator measurements can be executed fast and with high fidelity. These can be executed largely in parallel on non-overlapping stabilizers. Parallelization of flag-FT stabilizer measurements has been recently shown for distance-5 codes [20]. Also on stabilizers, which overlap on a number of data qubits, measurement circuits may be executed in parallel with appropriate gate schedules. Note that in a stabilizer-measurement-based state preparation routine, entangling gates are always only applied between data qubits and auxiliary qubits but never between two data qubits. As a consequence, the occurrence of dangerous mixed-type errors, as described in Sec. II.2.2, is naturally prevented for without additional flag qubits (apart from those that might be needed for FT stabilizer measurements, depending on the specific code).
II.4.2 Neutral atom platform
@@ II.4.3 Ion trap processor
III Numerical estimation of logical failure rates
Rate set | high () | low () |
In this section we investigate the logical failure rate of the fault-tolerant -gate implementation for the Steane code enabled via transversal code switching. The relatively low number of physical CNOT gates and its transversality properties play out in favor of achieving lower logical failure rates than the state-of-the-art flag-FT code switching scheme [41, 42] and even magic state injection [6, 52]. We do not include the intermediate error correction, described in Sec. II.2.1, in our numerical analysis.
@@ III Numerical estimation of logical failure rates
III.1 Depolarizing noise model
@@
The break-even point where can be identified at . This point is a meaningful comparison of an FT vs. physical-qubit implementation in the following sense: In a current or near-term quantum processor, a quantum algorithm performed with physical qubits will be dominated by the entangling gate error . For an implementation with 2D color codes, a quantum algorithm performed with logical qubits will be dominated by the logical -gate error since this is by far the most complicated logical gate to execute for this class of QEC codes. Note that this break-even point is very close to the actual code threshold of 2D color codes, , under circuit-level depolarizing noise [84, 85, 86], which reflects the low overhead of the code switching scheme and its advantageous design for operational noise.
Simulation of a distance-5 version of the protocol yields a lower logical failure rate than the distance-3 implementation for a physical error rate below . A more detailed discussion of the distance-5 simulation is postponed to Sec. V.
III.2 Multi-parameter noise model
@@
IV Comparison to other universality strategies
@@
In this section we provide a quantitative comparison of our transversal code switching scheme with other state-of-the-art techniques to achieve an FT universal gate set: Flag-based code switching in Sec. IV.1, magic state injection in Sec. IV.2 and concatenation in Sec. IV.3. All three techniques allow one to establish FT circuits on the physical level without employing complicated magic state distillation schemes, which are only FT when applied to logical qubits. A comparison of logical failure rates and resource count for specific small instances of these schemes is given in Tab. 2. Transversal code switching from the Steane code to the Tetrahedral code and back requires at least 24 qubits555It is assumed that the flag qubits are re-used after state preparation of either logical auxiliary qubit, otherwise the qubit count might increase to 27. Also, one may use a completely fresh set of qubits to prepare the Steane code auxiliary qubit in the second switching step, for instance if no fast reset operation is available. This would add another 7 data qubits and 1 flag qubit. and 83 CNOT gates.
@@
IV.1 Flag-style code switching

### v2

@@ Abstract
Code switching is an established technique that facilitates a universal set of fault-tolerant (FT) quantum gate operations by combining two quantum error correcting (QEC) codes with complementary sets of gates, which each by themselves are easy to implement fault-tolerantly. In this work, we present a code switching scheme that respects the constraints of FT circuit design by only making use of transversal gates. These gates are intrinsically FT without additional qubit overhead. We analyze application of the scheme to low-distance color codes, which are suitable for operation in existing quantum processors, for instance based on trapped ions or neutral atoms. We also briefly discuss connectivity constraints that arise for architectures based on superconducting qubits. Numerical simulations of circuit-level noise indicate that a logical -gate, facilitated by our scheme, could outperform both flag-FT magic state injection protocols and a physical -gate at low physical error rates. Transversal code switching naturally scales to code pairs of arbitrary code distance. We observe improved performance of a distance-5 protocol compared to both the distance-3 implementation and the physical gate for realistic error rates. We discuss how the scheme can be implemented with a large degree of parallelization, provided that logical auxiliary qubits can be prepared reliably enough. Our logical -gate circumvents the need for potentially costly magic state factories. The requirements to perform QEC and to achieve an FT universal gate set are then essentially the same: Prepare logical auxiliary qubits offline, execute transversal gates (ideally in parallel) and perform fast-enough measurements. Transversal code switching thus serves to enable more practical hardware realizations of FT universal quantum computation. The scheme alleviates resource requirements for experimental demonstrations of quantum algorithms run on logical qubits.
1 Introduction
@@
A prominent modern technique to render quantum circuits FT with little qubit and gate overhead is the flag qubit paradigm [12, 13, 14, 15, 16, 17, 18, 19, 20]. Faults within the circuit that may cause uncorrectable errors on the logical qubit are heralded by controlled propagation onto extra auxiliary qubits via additional CNOT gates, which are interleaved into the quantum circuit in an intricate way. When such a flag qubit is measured in a particular state, one may employ further circuity to treat such faults or repeat the circuit altogether until the flag is clear. On the one hand, flag circuits may be favorable due to their conceptual simplicity and typically small qubit overhead. On the other hand, dynamical circuit branching with relatively long circuit sequences might pose a challenge for practical applications of QEC with flag circuits. Also, recent developments of quantum computing hardware suggest that the number of qubits may no longer be the most critical limiting factor [21, 22, 23, 24, 25].
@@
For the controlled manipulation of encoded information, FT gate operations must be performed on the logical qubits. Transversal CNOT gates are appealing because they need not induce a time-overhead to respect fault tolerance such as, e.g., repeated operator measurements in lattice surgery [30, 31, 32]. It is known that no non-trivial QEC code offers a transversal universal gate set [33]. Therefore, given any code with a set of transversal gates, additional strategies are imperative to establish universality. There exist several procedures such as gate teleportation [34] or code conversion [35, 36] that vary in intricacy from both a theoretical and experimental point of view. Code switching in particular refers to the idea of employing two different QEC codes, which each on their own have a different set of transversal gates but in conjunction offer an FT universal gate set that is simple to perform on the logical level [37, 38]. The difficulty of implementing a particular gate is traded in for transferring the logical information between the two codes at will in an FT fashion.
@@
In this work, we investigate a new approach to implement the non-Clifford -gate fault-tolerantly on the 2D color code via FT code switching to the 3D color code that is based on code doubling [43, 44, 45, 46]. A notable strength of the scheme is that it can be viewed as deterministic in practice: Only QEC codes of distance are employed, instead of mere quantum error detecting codes with , and no post-selection is required. The scheme is considered efficient because it relies on a transversal construction of the logical CNOT gate, which works as long as one restricts the connection of the gate to be one-directional. Transversal code switching is performed via logical state teleportation. The logical -gate implementation is then built on two switching steps “2D 3D 2D”, depicted with logical building blocks in Fig. 1. Logical auxiliary qubits in both codes can be prepared with Clifford circuits only, so that no complicated magic state subroutines are needed. We find that transversal code switching can outperform other universality schemes in terms of logical failure rate, entangling gate count and circuit depth while only slightly increasing the number of required physical qubits.
This manuscript is organized as follows: The efficient FT -gate construction is discussed in detail in Sec. 2. We offer insights on parallelization and embedding into a reduced-connectivity hardware architecture. In Sec. 3, we present simulations of incoherent circuit-level Pauli noise with both a simple single-parameter depolarizing noise model and a more elaborate multi-parameter noise model that can be used to capture the essential noise processes of a trapped-ion quantum computer. A comparison to other strategies for establishing an FT universal gate set is provided in Sec. 4. In Sec. 5, we consider scaling up our scheme to QEC codes of arbitrary distance and lastly draw conclusions and provide an outlook on future work in Sec. 6.
2 Transversal code switching
@@
Essential for any distance- QEC code, FT code switching must ensure that no more than errors of any Pauli type or can result from up to fault events in a circuit-level noise model. Such faults must be assumed to happen during the execution of the actual quantum circuits of the switching algorithm where every physical operation acting on qubits is subjected to a -qubit depolarizing noise channel. A logical gate is deemed transversal if it can be implemented by applying bitwise physical gate operations to all physical qubits and, crucially, no interactions between physical qubits happen within the same logical block. Therefore, faults are confined to the qubits on which they initially occur and never spread uncontrollably within a code block.
@@
2.1 Doubled-code construction
@@
In Ref. [46], a code construction is described that allows one to perform FT code switching with the help of logical auxiliary qubits and one-way transversal CNOT gates. The central ingredients are a self-dual Calderbank-Shor-Steane (CSS) code111This need not be a color code. and a triorthogonal code, from which a so-called doubled code is constructed, which is itself again a triorthogonal code with parameters . The stabilizers and logical operators of the two ingredient codes are recombined to a new, bigger triorthogonal matrix in the process of code doubling. The initial self-dual code is complemented with additional -type stabilizers to yield the new triorthogonal code, which may have non-local stabilizers. Importantly, the original -type stabilizers of the self-dual code remain intact and are incorporated into the doubled code. As a consequence, the one-way logical CNOT gate can be executed by connecting physical CNOT gates in a pair-wise fashion between the physical qubits of the self-dual code and a subset of physical qubits of the doubled code. This enables bi-directional switching between the two codes via standard quantum state teleportation circuits at the logical level. Transversality ensures that fault tolerance is respected for a code switching step. Additionally, it must be guaranteed that no uncorrectable error can ever propagate between the two codes from the preparation of the logical auxiliary qubits. While code doubling has been investigated before [43, 44, 45], the one-way transversal CNOT gate is a new ingredient identified in Ref. [46].
2.2 Reduction to 2D/3D color code combination
@@
The logical CNOT gate can be performed directly between a 2D and a 3D color code at the price of restricting it to only one direction. The reason for this is that the 3D color code already is a doubled code. For the code definitions used in Ref. [46], this means that the control of the logical CNOT may only be encoded in the 3D color code and the target of the logical CNOT must be encoded in the 2D color code. It can be ensured that all stabilizer operators of each code propagate through the logical CNOT onto stabilizer operators of the respective other code, using only physical CNOT gates. The product of the stabilizer groups of both codes is therefore left invariant. Simultaneously, logical operators of both - and -type are mapped between the two codes as prescribed by the action of a logical CNOT gate.
@@
The remainder of this section is dedicated to applying the aforementioned code switching protocol to the smallest instances of the 2D and 3D color code respectively depicted in Fig. 2, i.e., the seven-qubit Steane code and the fifteen-qubit Tetrahedral code. Both QEC codes have distance and thus correct arbitrary Pauli error. FT state preparation of the logical auxiliary qubit states will be discussed in Sec. 2.2.2. After providing a detailed discussion of this small-scale implementation, we broaden the horizon again and discuss scaling up to large distances in Sec. 5.
2.2.1 Distance-3 transversal code switching
@@
We now also consider the propagation of errors. It is obvious that a single correctable error in the Steane code can, due to transversality, at most propagate to a single error on the Tetrahedral code, which is therein also correctable. Correctable -errors of weight-3 in the Tetrahedral code may either propagate to correctable weight-1 errors in the Steane code through a single physical CNOT gate but can also, and more severely, directly propagate to uncorrectable weight-3 -errors on the Steane code, as shown in Fig. 4b. Their appearance must therefore be prevented by an appropriate FT state preparation circuit. It is not enough to only fault-tolerantly prepare the 3D color code state but the state preparation routine must be FT with respect to the 2D color code as well in order to render the whole code switching scheme FT.
@@ 2.2.2 FT state preparation
2) High-weight -errors: -errors of weight larger than 1 on must be prevented from propagating to the Steane code during the course of code switching although they could be correctable in the Tetrahedral code. It is clear from Figs. 1 and 4b that propagation of such an error to the Steane code via the first logical CNOT gate, may cause a logically-flipped measurement result and subsequent erroneous application of a logical -operator. We find that, in our state preparation subroutine, three flags are necessary to catch all high-weight -errors. These flags can be chosen to measure the stabilizers , and .
@@
For practical purposes, we suggest to use these circuits for non-deterministic logical state preparation because the acceptance rate of approximately 333This value is estimated from a circuit-level depolarizing noise model with parameters and . More detail will be provided in Sec. 3. is reasonably close to unity for state-of-the-art quantum processors where the entangling gate error rate of order is the limiting factor [52]. No mid-circuit measurements or branching logic are required to correctly flag all dangerous faults. Deterministic variants of logical state preparation might as well be employed but we suspect this approach to require a larger number of entangling gates in general for low physical error rates.
@@ 2.3 Parallelized implementation
Given logical auxiliary qubits and parallelized physical-level gates, the transversal code switching scheme can be implemented in constant circuit depth w.r.t. the code size. The first logical CNOT requires one time step, the transversal measurements and corrections require one time step each, the physical -gates are all applied in parallel and the second code switching step then is done again in three time steps; the full logical -gate circuitry is thus performed in seven steps regardless of the code distances, assuming a supply of logical auxiliary qubits is available. Unitary encoding circuits for the preparation of -qubit logical auxiliary states with depth can be found for 2D topological codes with nearest-neighbor connectivity [59] and might be improved to by allowing for non-local qubit connectivity, exploiting symmetries [60] or with the help of physical auxiliary qubits [61].
@@ 2.4.1 Superconducting architecture
The long-range connectivity required for unitary encoding circuits could in theory be circumvented with the help of circuit knitting [64]. One may resort to a stabilizer-measurement-based state preparation routine altogether if non-destructive operator measurements can be executed fast and with high fidelity. These can be executed largely in parallel on non-overlapping stabilizers. Parallelization of flag-FT stabilizer measurements has been recently shown for distance-5 codes [20]. Also on stabilizers, which overlap on a number of data qubits, measurement circuits may be executed in parallel with appropriate gate schedules. Note that in a stabilizer-measurement-based state preparation routine, entangling gates are always only applied between data qubits and auxiliary qubits but never between two data qubits. As a consequence, the occurrence of dangerous mixed-type errors, as described in Sec. 2.2.2, is naturally prevented for without additional flag qubits (apart from those that might be needed for FT stabilizer measurements, depending on the specific code).
2.4.2 Neutral atom platform
@@ 2.4.3 Ion trap processor
3 Numerical estimation of logical failure rates
In this section we investigate the logical failure rate of the fault-tolerant -gate implementation for the Steane code enabled via transversal code switching. The relatively low number of physical CNOT gates and its transversality properties play out in favor of achieving lower logical failure rates than the state-of-the-art flag-FT code switching scheme [41, 42] and even magic state injection [6, 52]. We do not include the intermediate error correction, described in Sec. 2.2.1, in our numerical analysis.
@@ 3 Numerical estimation of logical failure rates
3.1 Depolarizing noise model
@@
The break-even point where can be identified at . This point is a meaningful comparison of an FT vs. physical-qubit implementation in the following sense: In a current or near-term quantum processor, a quantum algorithm performed with physical qubits will be dominated by the entangling gate error . For an implementation with 2D color codes, a quantum algorithm performed with logical qubits will be dominated by the logical -gate error since this is by far the most complicated logical gate to execute for this class of QEC codes. Note that this break-even point is very close to the threshold of QEC cycles with the 2D color code under circuit-level depolarizing noise, , reported in Refs. [84, 85, 86], which reflects the low overhead of the code switching scheme and its advantageous design for operational noise.
Simulation of a distance-5 version of the protocol yields a lower logical failure rate than the distance-3 implementation for a physical error rate below . A more detailed discussion of the distance-5 simulation is postponed to Sec. 5.
3.2 Multi-parameter noise model
@@
4 Comparison to other universality strategies
@@
In this section we provide a quantitative comparison of our transversal code switching scheme with other state-of-the-art techniques to achieve an FT universal gate set: Flag-based code switching in Sec. 4.1, magic state injection in Sec. 4.2 and concatenation in Sec. 4.3. All three techniques allow one to establish FT circuits on the physical level without employing complicated magic state distillation schemes, which are only FT when applied to logical qubits. A comparison of logical failure rates and resource count for specific small instances of these schemes is given in Tab. 2. Transversal code switching from the Steane code to the Tetrahedral code and back requires at least 24 qubits555It is assumed that flag qubits are re-used after state preparation of either logical auxiliary qubit, otherwise the qubit count might increase to 27. One may use a completely fresh set of qubits to prepare the Steane code auxiliary qubit in the second switching step if no fast reset operation is available. This would add another 7 data qubits and 1 flag qubit. and 83 CNOT gates.
@@
4.1 Flag-style code switching